(5.2.22) Assume that when adults with smartphones are randomly​ selected, 36​% use them in meetings or classes. If 25 adult smartphone users are randomly​ selected, find the probability that exactly 17 of them use their smartphones in meetings or classes.
The probability is? (round to four decimal)
1
Expert's answer
2020-12-01T01:44:15-0500

The 17 users from 25 can be chosen in "{25 \choose 17}" ways . Chance that 17 users from 25 use their smartphones in meetings or classes is "0.36^{17}*0.64^8" . The final probability is

"{25 \choose 17}*0.36^{17}*0.64^8=\frac {25!} {17!8!}*0.36^{17}*0.55^8=0.0003" or 0.03%